OpenStudy (anonymous):

18. Angela and Neil are going to the movies. They each bought a medium popcorn, and Neil got a small soft drink. Angela had a $5 gift certificate to put toward the cost, and Neil paid the rest, which came to $27.90. A movie ticket costs $10.00, and a medium popcorn costs $5.50. How much does a soft drink cost at the theater? (1 point) $1.90 $7.40 $2.90 $17.40

OpenStudy (anonymous):

Well you know that the total of the movie tickets is 20 dollars and the total of the medium popcorn is 11 dollars. The total was 27.90 which Neil paid plus the gift card 5 dollars so the equation is : x + 31 = 32.90 x= 32.90 - 31 x = 1.90 dollars the soft drink costs 1 dollar and 90 cents which is the first choice

OpenStudy (anonymous):

32.90-popcorn ($11) - tickets ($20) - gift certificate ($5) = $1.90